OpenStudy (anonymous):

Find the least common denominator of 1/4 and 11/6

OpenStudy (anonymous):

first find multiples of 4 then find multiples of 6 then find the lowest multiple out of the two

OpenStudy (anonymous):

4x6 = 24 thats the lcd

OpenStudy (anonymous):

Write the LCD for each pair of fractions. 13. 1/3,1/5 14. 2/7,1/4 15. 3/4,3/5 i dont get this question,i dont know how 2 find the lowest common denominator.please help me because this is going 2 be on a test in 2 dayz Hi, Adding two fractions with the same denominator is easy. 1/5 + 2/5 = (1 + 2)/5 = 3/5. If you have two fractions with different denominators, you need to find equivalent fractions with the same denominators and then add these fractions. So for example with your first problem 1/3 + 1/5 You can write 1/3 = 5/15 and 1/5 = 3/15 so 1/3 + 1/5 = 5/15 + 3/15 = 8/15. So, how did I know to use a denominator of 15? I used 15 because 15 is a multiple of both 3 and 5, and 15 is the smallest number that is a multiple of both 3 and 5. That is, 15 is the least common multiple (LCM) of 3 and 5. Likeise in your second problem. The denominators are 7 and 4 and 28 =7 4 is a multiple of both 7 and 4 so you can write 2/7 + 1/4 = 8/28 + 7/28 = 15/28 I changed 2/7 to 7/28 by multiplying both the numerator and denominator by 4, and I changed 1/4 to 7/28 by multiplying both the numerator and denominator by 7. Smetimes it is not that easy. If you are to add 3/4 and 1/6 then 4 6 = 24 is a multiple of 4 and 6 but there is a smaller multiple, 12. Thus you can write 3/4 + 1/6 = 9/12 + 2/12 = 11/12. In this case 12 is the least common multiple of 4 and 6 so 12 is the least common denominator of 3/4 and 1/6. To find the least common multiple of two numbers first find the prime factorization of the numbers 4 = 22 6 = 2 3 Look at the highest power of each of the primes that appear. In this case the hightes power of 2 is 2 and the highest power of 3 is 1. Form the number 22 3 = 12 This is the least common multiple of 4 and 6. Penny

OpenStudy (sweetburger):

omg thats a great explanation medal for that